Question: Find the equation of the line that passes through points (7,-2) and (5,1)
Write the equation of the line in slope-intercept form.
I know the the slope intercept for is y = mx +b

Answer by xcentaur(357)

Let the points be P(7,-2) and Q(5,1).


Slope of a line (m) =
= (difference of ordinates)/(difference of abcissae)
= (y2-y1)/(x2-x1)
= (1-(-2))/(5-7)
= (1+2)/(2)
= (3/2)


Now m=3/2


Taking form of:
y-y1=m(x-x1)


y-(-2)=3/2(x-7)
y+2=3/2(x-7)
2y+4=3(x-7)
2y+4=3x-21
2y-3x=-21-4
2y-3x=-25


Converting to standard form of slope-intercept we get,
2y=3x-25
Here,slope(m)=(3/2)
yintercept=(-25/2)


Hope this helps,
best of luck